Tina is training for a biathlon. As part of her training, she runs an 7-mile course, 4 miles of which is on level ground and 3 miles of which is downhill. It takes her 40 minutes to run the level part of the course and 20 minutes to run the downhill part of the course. Find her average speed in minutes per mile and in miles per minute for each part of the course. Round to the nearest hundredth, if rounding is necessary.

Answer by ankor@dixie-net.com(22740)

she runs an 7-mile course, 4 miles of which is on level ground and 3 miles of which is downhill.
It takes her 40 minutes to run the level part of the course and 20 minutes to
run the downhill part of the course.
Find her average speed in minutes per mile and in miles per minute for each part of the course.
:
The level part:
40/4 = 10 min/mi, and 4/10 = .4 mi/min
:
down hill part;
20/3 = 6.67 min/mi, and 3/6.67 = .45 mi/min